An oil refinery has decided to purchase some new drilling equipment for $550,000. The equipment will be kept for 10 years before being sold. The estimated salvage value (SV) for depreciation purposes is to be $25,000. Use this information to solve the following questions:

a) Using the straight line (SL) method, the annual depreciation on the equipment is _________________.

b) Using the double declining balance (DDB) method, the depreciation charge in year 3 is ______________.

c) Using the SL method, the book value (BV) at the end of the depreciable life is ____________________.

d) If SL depreciation is used and the equipment is sold for $35,000 at the end of the 10 years, the taxable gain (capital gain) on the disposal of the equipment is _________________.

e) If a MACRS depreciation schedule is used (based on a 5-year property class*) and the equipment is sold for $35,000 at the end of the 10 years, the taxable gain (capital gain) on the disposal of the equipment likely to be _________________. (Percentages: 20.00, 32.00, 19.20, 11.52, 11.52, 5.76)
Answer #1

(a)

SL annual depreciation ($) = (Cost - Salvage value) / Useful life = (550,000 - 25,000) / 10 = 525,000 / 10 = 52,500

(b)

SL depreciation rate = 1 / Useful life = 1/10 = 0.1

DDB depreciation rate = 2 x SL rate = 2 x 0.1 = 0.2

DDB depreciation, year 1 = $550,000 x 0.2 = $110,000

Book value at end of year 1 = $(550,000 - 110,000) = $440,000

DDB depreciation, year 2 = $440,000 x 0.2 = $88,000

Book value at end of year 2 = $(440,000 - 88,000) = $352,000

DDB depreciation, year 3 = $352,000 x 0.2 = $70,400

Book value at end of year 3 = $(352,000 - 70,400) = $281,600

(c)

When SL method is used, book value at end of year 10 (project life) = Salvage value = $25,000

(d)

Capital gain = Sale value - Book value = $35,000 - $25,000 = $10,000

(e)

Accumulated depreciation using MACRS = $550,000 x (20 + 32 + 19.2 + 11.52 + 11.52 + 5.76)% = $550,000 x 100%

= $550,000

Capital gain = Sale value - Book value = $35,000 - $0 = $35,000